OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope. C(0, 4), m = 0

OpenStudy (anonymous):

Plug in the formula: \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

so that would be 0-01+m(4-41)

OpenStudy (anonymous):

Forgot to mention, you only plug in the y1 and x1, like so: y - 4 = 0 (x - 0)

OpenStudy (anonymous):

Then you solve for y.

OpenStudy (anonymous):

and thats standard form?

OpenStudy (anonymous):

y=4 would be the standard form.